# PlasmoData.jl – A Julia Framework for Modeling and Analyzing Complex Data as Graphs

### v1

@@ 3.2.2 Euler Characteristic
The EC is a scalar integer value that summarizes the shape of a topological space [19, 67]. For a graph, this can be defined as [19]:
@@ 4 Case Studies
We now provide case studies that highlight different applications and benefits of the DataGraph abstraction. All scripts and data necessary to replicate the results in this section are available at https://github.com/zavalab/JuliaBox/tree/master/PlasmoData_examples.
@@ 4.1 Image Analysis
We consider a case study arising in image classification and show how the DataGraph abstraction and PlasmoData.jl help integrate with machine learning tools and facilitate feature extraction. Chemoresponsive liquid crystals (LCs) have garnered increased interest because of their potential as sensors for gas contaminants [70, 71, 72]. LCs have been shown to elicit different, visible chemical responses to varying chemical environments as the LCs change their orientation ordering [73, 74]. One of the challenges that naturally arises from these sensors is identifying the chemical environment surrounding a LC based on the appearance of the LC’s surface. We are therefore interested in being able to classify images of LCs based on the chemical environment into which the LC has been placed. We will show how these LC surfaces can be analyzed by representing images of the surfaces using the DataGraph abstraction, and we will show how we use PlasmoData.jl for extracting features from this data and use those features to train a machine learning model. We will also highlight how PlasmoData’s data structure can be employed for GNNs.
The dataset are images of LCs exposed to four different concentrations of sulfur dioxide (SO2): $0.5$, $1.0$, $2.0$, and $5.0$ ppm (i.e., four classes of images corresponding to each concentration). Each image is $134\times 134$ pixels, and there are 72 images in each class (288 total). Examples of images from each class can be seen in Figure 9[73, 69].
Figure 9: Images of LCs exposed to gaseous environments with different concentrations of SO2 [73, 69].
@@ 4.1.1 Classification using TDA and SVMs
Each of the 288 images can be treated as a tensor, $T\in[0,1]^{134\times 134\times 3}$. Under the DataGraph abstraction, we can represent the $i$th tensor as a graph with (16) using the first method shown in Section 2.3.2, where the third dimension of the tensor represents the red, green, and blue channels of the pixel intensity and where $t_{j,k,l}$ is the $j,k,l$ entry of $T$.
@@
The code for performing the tasks of generating the graph representations, building the EC curves, and training an SVM can be seen in Code Snippet 7. The data is in the format of a 4D array of size $(288,134,134,3)$. The graph is formed by treating each pixel of the image (dimensions 2 and 3 of the array) as a node using PlasmoData.jl’s matrix_to_graph function (Line 17). The intensity values for each RGB channel (dimension 4 of the array) are saved as weights on each node of the graph. After each graph is formed, we then find the EC curve for each weight (channel) on the graph and store this in the array so2_ECs (Lines 19 - 23). The three EC curves (one for each channel) are then concatenated into a vector and used to train a linear SVM using 5-fold cross validation (CV) (Lines 29 - 50).
@@
1        using PlasmoData, JLD, MLUtils, LIBSVM
@@
3        ### Construct matrices as graphs and get EC curves ###
@@
5        # Load in the SO2 data; so2_data is size (288, 134, 134, 3)
6        so2_data = load("so2_data.jld")["data"]
7        so2_classes = load("so2_classes.jld")["classes"]
@@
9        # Define threshold range for each EC curve
10        thresh = 0:.005:1
@@
12        # Define a matrix for the EC curves; EC curves will be concatenated
13        so2_ECs = Array{Float64, 2}(undef, (length(thresh)*3, 288))
@@
26        ### Perform 5 fold CV with SVMs on EC data ###
@@
28        # shuffle data
29        Xs, ys = shuffleobs((so2_ECs, so2_classes))
@@
31        # define a function for calculating accuracy
32        function get_accuracy(yhat, ytest)
@@
42        # Perform 5-fold CV
43        accuracy_values = []
44        for (train_data, val_data) in kfolds((Xs, ys); k = 5)
@@
46            yhat, decision_values = svmpredict(model, val_data[1])
47
@@
50        end
The method can effectively classify these images in a reasonable time. The classification accuracy, based on 5-fold CV, was $94.8\%$ with a standard deviation of $2.1\%$. Running this script on an Intel(R) Core i9-10885H (2.40GHz) processor resulted in a data processing time (Lines 10 - 24) of $36.7$ s and a 5-fold CV (Lines 29 - 50) time of $0.08$ s. This is comparable to what Jiang and co-workers[69] found, where they had $87.5\%$ accuracy from other topological methods that took less time to process and train ($\sim 14$ s). For comparison, the authors also trained a convolutional neural network (CNN) for classifying these images, and the CNN was able to classify with $95.1\%$ accuracy while taking $154$ seconds to train. Here, we do not aim to show superiority of one method over another, but to provide a general validation as to whether our methods yield reasonable results. In particular, our results suggest that it is possible to obtain high accuracy while performing these operations in a competitive time.
We also note—as Jiang and co-workers did in their study—that there are benefits to topological methods such as those we followed above. First, the linear SVM above had relatively few parameters—each EC curve contained $201$ points before concatenation, so there were only $603$ parameters in the SVM. In contrast, CNNs can use high numbers of parameters (thousands to millions). They also can be sensitive to rotation, whereas graphs are rotationally invariant (e.g., the EC curve does not depend on the orientation of the graph). In addition, Jiang and co-workers note that the CNN in their study was trained on a GPU, whereas our above methods (and Jiang and co-workers TDA methods) functioned efficiently on a CPU.
@@
Figure 10: Examples by class of graph filtrations at different threshold values. Each image has a value corresponding to the RGB channels, and these channels result in different topologies at the filtration thresholds. The EC curve captures these different topologies across channels and across classes. Data comes from the work of Bao et al. [73].
Figure 11: EC Curves for the RGB channels of four classes of images for LC data.
Because the EC of a graph is only the number of nodes minus the number of edges, the EC curve begins at the origin but can end far from the origin. In the case of images, the full graph (i.e., when nothing is filtered out) has a large negative EC since the number of edges is much more than the number of nodes. Thus, the EC curve of a graph tends to show when and how quickly edges and cycles begin to appear. For example, in the case of images, if the filtration results mostly in nodes that are far apart, there are few if any edges between them and the EC is more positive. In contrast, if the filtration results in groups of clustered nodes, there are more edges connecting nodes and the EC is more negative. This helps explain why the EC curves for the $0.5$ ppm SO2 class tend to be more positive at a given threshold value than the other classes (Figure 11) since, as can be seen in Figure 10, the filtration results in more spread out, small clusters at the lower threshold values for the $0.5$ ppm SO2 class than the other classes.
@@ 4.1.2 Classification using Graph Neural Networks
1        using PlasmoData, JLD, Flux, GraphNeuralNetworks
@@ 8            GNN_graphs = []
24        end
With the data formatted as GNNGraphs, we can build a full GNN through GraphNeuralNetworks.jl. We ran our script on an AMD EPYC 7302 16-core processor with access to a NVIDIA QUadro RTX 6000 GPU (used for training the GNN). Using the same 288 images as before, we were able to classify the images with $86.5\%$ accuracy with a standard deviation of $4.4\%$ based on 5-fold CV. While the time for data processing was short ($4.04$ s), the GNN did take a long time to train and perform CV (917 s).
This framework for building GNNs from PlasmoData.DataGraphs could be applied much more broadly than just to images (e.g., molecular property prediction). The user has significant flexibility in constructing their own graph data and using GNNs with that data. For further details on constructing a GNNGraph from data using PlasmoData.jl, see the Supporting Information.
@@ 4.2 Multivariate Time Series Analysis
In this example, we illustrate how multivariate time series can be represented as graphs, and we highlight how the topology of this graph representation can provide interesting insights.
@@
Recently, de Souza and co-workers [13] explored the use of TDA (and in particular the EC) as a potential tool for data-driven surveillance of epidemic outbreaks. Specifically, they were interested in finding indicators or fingerprints of when an outbreak is occurring (or likely to occur), and they based their efforts on locally reported disease data and on simulated data. In their work, they constructed correlation matrices as graphs, filtered these graphs, and looked at the resulting topology. They identified that there is a strong correlation between the EC curve and epidemic outbreaks, and they suggest that TDA could be used as a tool in disease surveillance. In this section, we will recreate some of their results using the DataGraph abstraction and using PlasmoData.jl, and we will suggest some other topological metrics that could be useful in analyzing datasets like these. In doing this, we emphasize that our purpose in this example is to highlight how data can be modeled with graphs and how PlasmoData.jl facilitates much of this analysis.
@@
First, we highlight the general data and methodology of de Souza and co-workers, but point the reader to their original paper to find an in depth explanation [13]. For this analysis, we will use their data from Recife, Brazil of new daily dengue cases from 93 individual districts from 2014 - 2021 (Figure 12). For each seven-day moving window of data, they computed the Pearson correlation for the 93 time series (one for each district) and then constructed the edge-weighted graph as outlined earlier (see for example Figure 6). They then computed a characteristic threshold value (what they call a critical percolation value) for this edge-weighted graph. This characteristic threshold value is the maximum possible threshold value that does not change the number of connected components from the number in the original graph when all edges with weights less than that threshold value are removed. A visualization of a this process can be seen in Figure 13, where $\ell$ is the threshold value used for filtration.
@@
Figure 12: New dengue cases by week for 93 districts in Recife, Brazil.
@@
Figure 13: General procedure used by de Souza and co-workers [13]. For a 7-day moving time window, they formed a (Pearson) correlation matrix and then constructed an edge weighted graph. They then filtered out all edges and progressively added edges back into the graph by changing the filtration threshold, $\ell$, until the number of connected components was equal to that of the original graph.
@@
Note that, in computing the EC of the filtered graph, de Souza and co-workers use methods different from that discussed in this present work. Where we have only discussed computing the EC for a graph (a 2D object), de Souza and co-workers treat the filtered graph as a CW-complex. CW-complexes are different than graphs, and involve the union of all d-cells in a graph (equivalently, the union of all $d$-node or -vertex cliques; see [13, 78]). A $d$-cell or $d$-vertex clique is a subset of nodes in a graph such that every node is connected by an edge to every other node in the subset (i.e., every two nodes in the subset are adjacent). A detailed discussion of CW-complexes is outside the scope of this paper; however, we mention this to highlight that the EC computed by de Souza and co-workers for CW-complexes is different than the EC we will show computed for a graph. Instead, their methods for computing the EC involve finding all $d$-cells up to some $d_{max}$ value. We will discuss later in this subsection some of the implications of this difference.
@@
We can represent the Pearson correlation matrix at each time point within the DataGraph abstraction, and we then can perform a similar analysis as de Souza and co-workers on the DataGraph within PlasmoData.jl. For the time series of data $\bm{X}\in\mathbb{Z}_{+}^{2922\times 93}$ and $\bm{X}=[X_{1},X_{2},...,X_{2922}]$, we first construct the Pearson correlation matrix $R_{t}=corr([X_{t-6},X_{t-5},...,X_{t}])$ for each time $t$ between 7 and 2922. Since $R_{t}$ is a symmetric matrix, we can construct a DataGraph following the method outlined in Section 2.3.3. The DataGraph for $R_{t}$ is given in (17), where $r_{i,j}$ is the $i,j$ entry of $R_{t}$.
@@
We performed the above analysis in PlasmoData.jl, and we will show how the DataGraph abstraction can enable topological analysis of the dengue data. PlasmoData.jl readily facilitates the process of building the graph, filtering the edges, and performing TDA. Code Snippet 9 shows how these tasks can be implemented. First, the data for Recife, Brazil from de Souza and co-workers is loaded on line 4. This data are daily reported new cases of dengue outbreaks in each of 93 city districts. We then define the function find_smallest_filtered_graph on Lines 7 - 25 that will add edges to the graph until the graph has only one connected componentent. The symmetric matrix for building the graph is formed by computing the Pearson correlation matrix on Line 34. Because we use a limited time window (7 days), some values in the matrix were not defined (i.e., NaN values). We removed all rows/columns of the symmetric matrix that were NaN values (Lines 37 - 38), and used the resulting symmetric matrix to construct the edge weighted graph, $\mathcal{DG}$ (line 46) (this results in a symmetric matrix with fewer than 93 nodes). We next define a sequence of values $\ell_{1},\ell_{2},...,\ell_{k}$ which are equal to the edge weights of the graph ($\bm{d}_{a}^{\mathcal{E}}$) in descending order (see Line 10; non-unique values are removed in the code). Once $\mathcal{DG}_{t}$ is defined, we then iterate through the sequnce of edges weights until $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=1$, where $\ell_{i}$ is the $i$th value in the sequence, and where $\ell_{i}=\epsilon_{t}$. Note that we use $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=1$ rather than $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=|\mathcal{DG}_{t}|$ because we have removed the rows and columns of the matrix which are not defined; this results in the edge-weighted PlasmoData.DataGraph formed in line 15 having only one connected component.
Code Snippet 9: Code for converting time series data to an edge-weighted graph. The code follows the general procedure of de Souza and co-workers [13] in filtering the graph.
1        using PlasmoData, Graphs, DelimitedFiles
@@
3        # Read in data
4        data = readdlm("Recife_data.csv", ’,’, Int)
@@
6        # Define function for performing filtration
7        function find_smallest_filtered_graph(graph)

### v2

@@ 3.2.2 Euler Characteristic
The EC is a valuable descriptor for topological objects, including graphs. The EC is a scalar integer value that summarizes the shape of a topological space [19, 67]. For a graph, this can be defined as [19]:
@@ 4 Case Studies
We now provide case studies that highlight different applications and benefits of the DataGraph abstraction. All scripts and data necessary to replicate the results in this section are available at https://github.com/zavalab/JuliaBox/tree/master/PlasmoData_examples. These examples include image analysis (representing matrices and tensors as graphs), multivariate time series analysis (representing symmetric matrices as graphs), and connectivity analysis (analyzing pathways within graphs).
@@ 4.1 Image Analysis
We consider a case study arising in image classification and show how the DataGraph abstraction and PlasmoData.jl help integrate with machine learning tools and facilitate feature extraction. We will show how these images of surfaces can be analyzed by representing these images as 3D tensors the representing these tensors using the DataGraph abstraction. We will show how we use PlasmoData.jl for extracting features from this data and use those features to train a machine learning model. We will also highlight how PlasmoData’s data structure can be employed for GNNs.
The images we consider are of chemoresponsive liquid crystals (LCs) which have garnered increased interest because of their potential as sensors for gas contaminants [70, 71, 72]. LCs have been shown to elicit different, visible chemical responses to varying chemical environments as the LCs change their orientation ordering [73, 74]. One of the challenges that naturally arises from these sensors is identifying the chemical environment surrounding a LC based on the appearance of the LC’s surface. We are therefore interested in being able to classify images of LCs based on the chemical environment into which the LC has been placed. The dataset we consider are images of LCs exposed to four different concentrations of sulfur dioxide (SO2): $0.5$, $1.0$, $2.0$, and $5.0$ ppm (i.e., four classes of images corresponding to each concentration). Each image is $134\times 134$ pixels, and there are 72 images in each class (288 total). Examples of images from each class can be seen in Figure 9 [69, 73].
Figure 9: Images of LCs exposed to gaseous environments with different concentrations of SO2 [69, 73]
@@ 4.1.1 Classification using TDA and SVMs
To classify these images, we will represent these images as tensors and then each tensor as a graph. Each of the 288 images can be treated as a tensor, $T\in[0,1]^{134\times 134\times 3}$. Under the DataGraph abstraction, we can represent the $i$th tensor as a graph with (16) using the first method shown in Section 2.3.2, where the third dimension of the tensor represents the red, green, and blue channels of the pixel intensity and where $t_{j,k,l}$ is the $j,k,l$ entry of $T$.
@@
The code for performing the tasks of generating the graph representations, building the EC curves, and training an SVM can be seen in Code Snippet 7. The data is in the format of a 4D array of size $(288,134,134,3)$. The graph is formed by treating each pixel of the image (dimensions 2 and 3 of the array) as a node using PlasmoData.jl’s matrix_to_graph function (Line 17). The intensity values for each RGB channel (dimension 4 of the array) are saved as weights on each node of the graph. After each graph is formed, we then find the EC curve for each weight (channel) on the graph and store this in the array so2_ECs (Lines 19 - 23). The three EC curves (one for each channel) are then concatenated into a vector and used to train a linear SVM using 5-fold cross validation (CV) (Lines 29 - 49).
@@
1using PlasmoData, JLD, MLUtils, LIBSVM
@@
3### Construct matrices as graphs and get EC curves ###
@@
5# Load in the SO2 data; so2_data is size (288, 134, 134, 3)
6so2_data = load("so2_data.jld")["data"]
7so2_classes = load("so2_classes.jld")["classes"]
@@
9# Define threshold range for each EC curve
10thresh = 0:.005:1
@@
12# Define a matrix for the EC curves; EC curves will be concatenated
13so2_ECs = Array{Float64, 2}(undef, (length(thresh)*3, 288))
@@
26### Perform 5 fold CV with SVMs on EC data ###
@@
28# shuffle data
29Xs, ys = shuffleobs((so2_ECs, so2_classes))
@@
31# define a function for calculating accuracy
32function get_accuracy(yhat, ytest)
@@
42# Perform 5-fold CV
43accuracy_values = []
44for (train_data, val_data) in kfolds((Xs, ys); k = 5)
@@
46 yhat, decision_values = svmpredict(model, val_data[1]
@@
49end
The method can effectively classify these images in a reasonable time. The classification accuracy, based on 5-fold CV, was $94.8\%$ with a standard deviation of $2.1\%$. Running this script on an Intel(R) Core i9-10885H (2.40GHz) processor on a single thread with Julia 1.7.3 resulted in a data processing time (Lines 10 - 24) of $36.2$ s and a 5-fold CV (Lines 29 - 49) time of $0.09$ s. This is comparable to what Jiang and co-workers [69] found, where they had $87.5\%$ accuracy from other topological methods that took less time to process and train ($\sim 14$ s). For comparison, Jiang and co-workers [69] also trained a convolutional neural network (CNN) for classifying these images, and the CNN was able to classify with $95.1\%$ accuracy while taking $154$ seconds to train. Here, we do not aim to show superiority of one method over another or provide a rigorous comparison in processing time but rather to provide a general validation as to whether our methods yield reasonable results. In particular, our results suggest that it is possible to obtain high accuracy while performing these operations in a competitive time.
We also note—as Jiang and co-workers [69] did in their study—that there are benefits to topological methods such as those we followed above. First, the linear SVM above had relatively few parameters—each EC curve contained $201$ points before concatenation, so there were only $603$ parameters in the SVM. In contrast, CNNs can use high numbers of parameters (thousands to millions). They also can be sensitive to rotation, whereas graphs are rotationally invariant (e.g., the EC curve does not depend on the orientation of the graph). In addition, Jiang and co-workers [69] note that the CNN in their study was trained on a GPU, whereas our above methods (and Jiang and co-workers [69] TDA methods) functioned efficiently on a CPU.
@@
Figure 10: Examples by class of graph filtrations at different threshold values. Each image has a value corresponding to the RGB channels, and these channels result in different topologies at the filtration thresholds. The EC curve captures these different topologies across channels and across classes. Data comes from the work of Bao and co-workers [73] and Jiang and co-workers [69]
Figure 11: EC Curves for the RGB channels of four classes of images for LC data from Bao and co-workers [73] and Jiang and co-workers [69]
The EC curve for each graph (and each channel) could provide insight into the image. Because the EC of a graph is only the number of nodes minus the number of edges, the EC curve begins at the origin but can end far from the origin. In the case of images, the full graph (i.e., when nothing is filtered out) has a large negative EC since the number of edges is much more than the number of nodes. Thus, the EC curve of a graph tends to show when and how quickly edges and cycles begin to appear. For example, in the case of images, if the filtration results mostly in nodes that are far apart, there are few if any edges between them and the EC is more positive. In contrast, if the filtration results in groups of clustered nodes, there are more edges connecting nodes and the EC is more negative. This helps explain why the EC curves for the $0.5$ ppm SO2 class tend to be more positive at a given threshold value than the other classes (Figure 11) since, as can be seen in Figure 10, the filtration results in more spread out, small clusters at the lower threshold values for the $0.5$ ppm SO2 class than the other classes.
@@ 4.1.2 Classification using Graph Neural Networks
1using PlasmoData, JLD, Flux, GraphNeuralNetworks
@@ 8 GNN_graphs = []
24end
With the data formatted as GNNGraphs, we can build a full GNN through GraphNeuralNetworks.jl. We ran our script on an AMD EPYC 7302 16-core processor with access to a NVIDIA Quadro RTX 6000 GPU (used for training the GNN). Using the same 288 images as before, we were able to classify the images with $89.2\%$ accuracy with a standard deviation of $5.1\%$ based on 5-fold CV. While the time for data processing was short ($3.8$ s), the GNN did take a long time to train and perform CV (947 s).
The ability to use PlasmoData.DataGraphs for GNNs is a powerful capability. This framework for building GNNs from PlasmoData.DataGraphs could be applied much more broadly than just to images (e.g., molecular property prediction). The user has significant flexibility in constructing their own graph data and using GNNs with that data, enabling application of PlasmoData.jl to a variety of fields or use cases. For further details on constructing a GNNGraph from data using PlasmoData.jl, see the Supporting Information.
@@ 4.2 Multivariate Time Series Analysis
In this example, we illustrate how multivariate time series can be represented as graphs, and we highlight how the topology of this graph representation can provide interesting insights. Recently, de Souza and co-workers [13] explored the use of TDA (and in particular the EC) as a potential tool for data-driven surveillance of epidemic outbreaks. Specifically, they were interested in finding indicators or fingerprints of when an outbreak is occurring (or likely to occur), and they based their efforts on locally reported disease data and on simulated data. In their work, they constructed correlation matrices as graphs, filtered these graphs, and looked at the resulting topology. They identified that there is a strong correlation between the EC curve and epidemic outbreaks, and they suggest that TDA could be used as a tool in disease surveillance. In this section, we will recreate some of their results using the DataGraph abstraction and using PlasmoData.jl, and we will suggest some other topological metrics that could be useful in analyzing datasets like these. In doing this, we emphasize that our purpose in this example is to highlight how data can be modeled with graphs and how PlasmoData.jl facilitates much of this analysis.
@@
First, we highlight the general data and methodology of de Souza and co-workers [13], but point the reader to their original paper to find an in depth explanation [13]. For this analysis, we will use their data from Recife, Brazil of new daily dengue cases from 93 individual districts from 2014 - 2021 (Figure 12). For each seven-day moving window of data, they computed the Pearson correlation for the 93 time series (one for each district) and then constructed the edge-weighted graph as outlined earlier (see for example Figure 6). They then computed a characteristic threshold value (what they call a critical percolation value) for this edge-weighted graph. This characteristic threshold value is the maximum possible threshold value that does not change the number of connected components from the number in the original graph when all edges with weights less than that threshold value are removed. A visualization of a this process can be seen in Figure 13, where $\ell$ is the threshold value used for filtration.
@@
Figure 12: New dengue cases by week for 93 districts in Recife, Brazil based on data from de Souza and co-workers [13]
@@
Figure 13: General procedure used by de Souza and co-workers [13]. For a 7-day moving time window, they formed a (Pearson) correlation matrix and then constructed an edge weighted graph. They then filtered out all edges and progressively added edges back into the graph by changing the filtration threshold, $\ell$, until the number of connected components was equal to that of the original graph
@@
Note that, in computing the EC of the filtered graph, de Souza and co-workers use methods different from that discussed in this present work. Where we have only discussed computing the EC for a graph (a 2D object), de Souza and co-workers [13] treat the filtered graph as a CW-complex. CW-complexes are different than graphs, and involve the union of all d-cells in a graph (equivalently, the union of all $d$-node or -vertex cliques; see [13, 78]). A $d$-cell or $d$-vertex clique is a subset of nodes in a graph such that every node is connected by an edge to every other node in the subset (i.e., every two nodes in the subset are adjacent). A detailed discussion of CW-complexes is outside the scope of this paper; however, we mention this to highlight that the EC computed by de Souza and co-workers for CW-complexes is different than the EC we will show computed for a graph. Instead, their methods for computing the EC involve finding all $d$-cells up to some $d_{max}$ value. We will discuss later in this subsection some of the implications of this difference.
@@
We can represent the Pearson correlation matrix at each time point within the DataGraph abstraction, and we then can perform a similar analysis as de Souza and co-workers [13] on the DataGraph within PlasmoData.jl. For the time series of data $\bm{X}\in\mathbb{Z}_{+}^{2922\times 93}$ and $\bm{X}=[X_{1},X_{2},...,X_{2922}]$, we first construct the Pearson correlation matrix $R_{t}=corr([X_{t-6},X_{t-5},...,X_{t}])$ for each time $t$ between 7 and 2922. Since $R_{t}$ is a symmetric matrix, we can construct a DataGraph following the method outlined in Section 2.3.3. The DataGraph for $R_{t}$ is given in (17), where $r_{i,j}$ is the $i,j$ entry of $R_{t}$.
@@
We performed the above analysis in PlasmoData.jl, and we will show how the DataGraph abstraction can enable topological analysis of the dengue data. PlasmoData.jl readily facilitates the process of building the graph, filtering the edges, and performing TDA. Code Snippet 9 shows how these tasks can be implemented. First, the data for Recife, Brazil from de Souza and co-workers [13] is loaded on line 4. This data are daily reported new cases of dengue outbreaks in each of 93 city districts. We then define the function find_smallest_filtered_graph on Lines 7 - 25 that will add edges to the graph until the graph has only one connected componentent. The symmetric matrix for building the graph is formed by computing the Pearson correlation matrix on Line 34. Because we use a limited time window (7 days), some values in the matrix were not defined (i.e., NaN values). We removed all rows/columns of the symmetric matrix that were NaN values (Lines 37 - 38), and used the resulting symmetric matrix to construct the edge weighted graph, $\mathcal{DG}$ (line 46) (this results in a symmetric matrix with fewer than 93 nodes). We next define a sequence of values $\ell_{1},\ell_{2},...,\ell_{k}$ which are equal to the edge weights of the graph ($\bm{d}_{a}^{\mathcal{E}}$) in descending order (see Line 10; non-unique values are removed in the code). Once $\mathcal{DG}_{t}$ is defined, we then iterate through the sequnce of edges weights until $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=1$, where $\ell_{i}$ is the $i$th value in the sequence, and where $\ell_{i}=\epsilon_{t}$. Note that we use $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=1$ rather than $|f_{\mathcal{N}}(\mathcal{DG}_{t},\mathcal{L}_{t,\ell_{i}})|=|\mathcal{DG}_{t}|$ because we have removed the rows and columns of the matrix which are not defined; this results in the edge-weighted PlasmoData.DataGraph formed in line 15 having only one connected component.
Code Snippet 9: Code for converting time series data to an edge-weighted graph. The code follows the general procedure of de Souza and co-workers [13] in filtering the graph
1using PlasmoData, Graphs, DelimitedFiles
@@
3# Read in data
4data = readdlm("Recife_data.csv", ’,’, Int)
@@
6# Define function for performing filtration
7function find_smallest_filtered_graph(graph)